An air bubble at the bottom of a lake 38.5 m deep has a volume of 1.00 cm3. If the temperature at the bottom is 4.7 ∘C and at the top 19.3 ∘C, what is the radius of the bubble just before it reaches the surface?
Answer #1

here

at the bottom of late for the pressure

P2 = 1.01 *10^5 + 1000 * 9.8 * 38.5

P2 = 478300 Pa

for the radius of bubble

as P1 * V1/T1 = P2 * V2/T2

478300 * 1/(4.7 + 273) = 1.01 *10^5 * 4pi/3 * r^3/(19.3 + 273)

solving for r

r = 1.06 cm

the radius of the bubble is 1.06 cm